Explain how fuel depletion and transmutation affect reactor operation and control.

Please give me URL if you used
Answer #1

In a nuclear reactor, every fission initiates another fission. For every nucleus undergoing fission, the minimum possible condition is that it must produce atleast one neutron that causes fission. And hence, the number of fissions occuring per unit time in a system must be constant.

The infinite multiplication factor can be expressed mathematically as,

k_{\infty } = ( neutron production from fission in one neutron generation) / ( neutron absorption in the foregoing neutron generation)

When a reactor is operated, the amount of fissile material present in the fuel assembly constantly decreases, owing to this, the infinite multiplication factor k_{\infty } decreases. Wherever the power is greatest, the decrease of k_{\infty } will take place. This causes the shift in power during cycle depletion. Hence the maintainance of power distribution has to be done effeciently.

The process of conversion of an isotope or a chemical element into another is termed as transmutation. It can also be explained as the ability to produce a particular type of nuclide which can be related to a neutron absorption and decay process.

Because of the elements produced by transmutation, the materials used in the reactors are subjected to damage, due to its reaction with the materials.This results in tthe change of reactors mechanical and physical properties of the materials by which it is produced.